A flywheel has a mass of 5.0 kg, inner radius of 8.00 cm, and outer radius of 14.2 cm. Determine

a) angular frequency to store 400 kJ of rotational energy

b) the average power and torque delivered by the flywheel if 400kJ can be stored in 6.67 s
Answer #1

Moment of inetia of flywheel is given by

I = \frac{1}{2}M(a^2 + b^2)

M = 5 kg

a = 8 cm = 0.08 m

b = 14.2 cm = 0.142 m

I = \frac{1}{2}*5*(0.08^2 + 0.142^2)

I = 0.06641 kg.m^2

a) Rotational KE is given by formula

KE = \frac{1}{2}Iw^2

400*10^3 = \frac{1}{2}*0.06641*w^2

w = 3.47*10^3 rad/s

b) Average power delivered is given by

Power = \frac{Energy}{Time}

Power = \frac{400*10^3}{6.67}

Power = 5.99*10^4 Watt

Now to find torque we can use equation below

Power = Torque * AngularFrequency

Power = \tau * \omega

5.99 * 10^4 = \tau * 3.47*10^3

\tau = 17.3 N.m